### v1

# Towards Automated Semantic Interpretability in Reinforcement Learning via Vision-Language Models

Abstract
Semantic Interpretability in Reinforcement Learning (RL) enables transparency, accountability, and safer deployment by making the agent’s decisions understandable and verifiable. Achieving this, however, requires a feature space composed of human-understandable concepts, which traditionally rely on human specification and fail to generalize to unseen environments. In this work, we introduce Semantically Interpretable Reinforcement Learning with Vision-Language Models Empowered Automation (SILVA), an automated framework that leverages pre-trained vision-language models (VLM) for semantic feature extraction and interpretable tree-based models for policy optimization. SILVA first queries a VLM to identify relevant semantic features for an unseen environment, then extracts these features from the environment. Finally, it trains an Interpretable Control Tree via RL, mapping the extracted features to actions in a transparent and interpretable manner. To address the computational inefficiency of extracting features directly with VLMs, we develop a feature extraction pipeline that generates a dataset for training a lightweight convolutional network, which is subsequently used during RL. By leveraging VLMs to automate tree-based RL, SILVA removes the reliance on human annotation previously required by interpretable models while also overcoming the inability of VLMs alone to generate valid robot policies, enabling semantically interpretable reinforcement learning without human-in-the-loop.
I Introduction
Reinforcement learning (RL) agents operate in dynamic environments, often developing policies that rely on latent representations not directly interpretable by humans [1]. Ensuring that these learned representations align with human-understandable concepts, a goal often referred to as semantic interpretability, is crucial for ensuring safety and preventing unintended harmful actions. While progress has been made in developing interpretable AI models, achieving semantic interpretability in RL remains particularly challenging  [2].
Semantic interpretability refers to the extent to which a model’s features, decisions, or internal representations can be understood and explained in terms of human-understandable concepts. Achieving this requires two essential components: (1) the inputs to the decision-making model must be intelligible to facilitate downstream interpretation, and (2) the decision-making process must be built with an interpretable model that allows humans to understand and explain its rationale  [3]. In essence, achieving semantic interpretability involves mapping high-dimensional input data to human-understandable features that can serve as a low-dimensional input space for interpretable decision-making models.
To learn an interpretable model on a semantically interpretable feature space, various model classes can be considered, including decision trees and rule-based policies, both of which facilitate explicit reasoning about an agent’s actions. Decision trees have been used to represent policies in a structured, human-readable format [4, 5, 6]. Similarly, rule-based methods, such as logical programs and symbolic reasoning approaches, offer an alternative by encoding policies in a form that can be explicitly followed and verified [7, 8]. Additionally, hierarchical decision-making frameworks have been explored, where high-level policies generate interpretable subgoals while low-level controllers handle execution [9, 10]. However, none of these prior works address high-dimensional image or video inputs; instead, they rely either on human-defined semantic concepts or ground-truth feature vectors provided by the environment.
Existing approaches have attempted to bridge the gap between raw image inputs and interpretable features [11] to facilitate model understanding but fall short in semantic interpretability. Many approaches either produce mathematical or visual representations that do not correspond to human-understandable concepts [12, 13] or rely on predefined feature spaces that aren’t generalize across different environments, using CNNs [14] or ViTs [15] to establish mappings between inputs and semantic features. As a result, human intervention has traditionally played a key role, either by manually defining semantically meaningful feature spaces [16] or by labeling machine-generated features to associate them with semantically meaningful concepts [17].
The emergence of Vision-Language Models (VLMs) has opened new possibilities for overcoming these challenges. These models, known for their capabilities in fine-grained visual understanding, language reasoning, and decision-making  [18, 19, 20], have seen widespread use in various tasks  [19, 21]. In this work, we demonstrate that VLMs can substitute humans in defining and extracting semantic features, serving not merely as a “scarecrow in oz”  [22] or a cost-effective fallback to human labor, but as a scalable and automated solution that matches human performance.
We propose Semantically Interpretable Reinforcement Learning with Vision-Language Models Empowered Automation (SILVA), which is to the best of our knowledge the first automated framework that achieves semantic interpretability in RL and generalizes effectively to unseen environments. By leveraging VLMs to extract features from raw environment inputs and training with Interpretable Control Trees, we manage to provide the rationale behind a RL algorithm in terms of human-understandable concepts. Notably, our framework (i) maintains competitive performance with black-box models while surpassing uninterpretable models built on the same feature space, (ii) supports both image and video inputs, effectively capturing temporal information, (iii) generalizes to unseen environments without human-in-the-loop, and (iv) enables computational efficiency by training a lightweight convolutional network as an alternative to VLM-based feature extraction.
II Related Work
II-A Interpretable Reinforcement Learning
Figure 1: Overview of SILVA, our proposed framework. A. Feature Extraction involves a three-step process where: (1) The VLM identifies key features given domain knowledge and environment observation. (2) The VLM generates code to extract features from a single image frame and assesses whether certain key features require multiple frames. If so, it builds upon the single-frame code to extract these features. (3) Input-output pairs are collected using the extractor code and are used to train an Extractor Convolutional Neural Network (Extractor CNN) as a computationally efficient substitute. B. Tree-based RL Training leverages the Extractor CNN to obtain semantic features from image observations, which are subsequently used as inputs to the policy network (an Interpretable Control Tree) to generate control actions for the RL agent.
Interpretable RL is inherently challenging compared to standard supervised learning, as it involves long-term consequences to actions and latent receptions of data through exploration.  [2] A significant line of research has explored interpretability in RL through generating visual saliency maps using Jacobian-based, perturbation-based, and attribution-based methods [23, 13, 24]. However, these approaches primarily focus on post-hoc explanations and fall out of a stricter definition of Interpretable RL  [3], which requires all components including the input space and the decision-making process to be intelligible. Approaches to achieve intelligible decision-making includes interpretable tree-structures, rule and decision lists, and other rule-based methods [3], among which decision trees are arguably the easiest to verify  [25]. Among state-of-the-art tree methods, Differentiable Decision Trees (DDTs) employ a Decision Tree structure that allows for online learning via stochastic gradient descent [25], while Interpretable Continuous Control Trees (ICCTs) extend this capability to continuous control tasks, ensuring both high performance and interpretability [26]. Similarly, POETREE (Policy Extraction through Decision Trees) [27] introduces dynamically growing probabilistic decision trees that adapt throughout training, allowing the tree complexity to evolve based on the task requirements. Our work addresses the limitation of prior tree-based methods by automating feature definition and extraction, thus enabling generalization to unseen environments.
II-B Feature Extraction for Reinforcement Learning
Common techniques that map high-dimensional inputs into low-dimensional feature spaces for RL includes linear nonparametric methods, such as manifold learning and spectral learning, and nonlinear parametric methods, such as deep learning  [12]. Other work focuses on the extraction of visual features that are not inherently connected with semantic meanings.  [23] employs a self-supervised interpretable network (SSINet) to produce fine-grained attention masks for highlighting task-relevant information.  [13] computed perturbation-based saliency of the image input to deduce regions that had the most effect. However, only a limited number of studies have proposed approaches without human-in-the-loop to construct semantic feature spaces for RL training.  [28] leverages information from Atari instruction manuals to construct the feature space, then uses unsupervised object detection and grounding to acquire the corresponding
objects in the environment. However, they observe that current techniques exhibit a lack of reliability  [28]. Our framework distinguishes from prior work by enabling automatic feature definition and extraction without human-in-the-loop.
II-C VLMs for Reinforcement Learning
VLMs have been applied in various aspects of RL training, including reward design, preference feedback, and representation deriving [29, 30, 20]. One prominent approach involves using VLMs to generate reward functions.  [29] use VLMs as zero-shot reward models to specify tasks via natural language, while  [31] leverages VLMs to produce dense reward functions through code generation. Additionally,  [30, 32] use VLMs to provide preference feedback with which they updated the reward functions for preference-based RL. Another line of work leverages VLMs as feature extractors to obtain valuable information which facilitates training.  [20] extract promptable representations from VLMs and decode output tokens into embeddings, which are used to initialize RL policies. However, those representations are not derived for interpretability purposes and might not often be correct. To the best of our knowledge, our work is the first to utilize VLM for creating a semantic feature space for RL training.
III Methodology
We introduce Semantically Interpretable Reinforcement Learning with Vision-Language Models Empowered Automation (SILVA), a framework designed to automatically learn an interpretable decision-making RL policy from an unseen environment. SILVA leverages the visual comprehension and reasoning capabilities of VLMs automate feature extraction through a three-step pipeline: (1) it employs zero-shot prompting on the environment to identify a set of semantic features that deterministically contribute to the decision-making, (2) it generates feature extraction codes that captures both static and temporal information, and (3) it collects a dataset with VLM-generated extraction codes to train a lightweight convolutional neural network (Extractor CNN) as a computationally efficient substitute. Subsequently, utilizing the Extractor CNN, SILVA performs tree-based RL training with an Interpretable Control Tree, establishing a transparent mapping between the semantic feature space and action outputs to derive a semantically interpretable policy for decision-making.
III-A Feature Extraction
The feature extraction process of SILVA involves a three-step pipeline, as shown in Fig. 1. In Step 1, we query the VLM to generate key features that contribute to the decision in the environment. In Step 2, we query the VLM to generate code that extracts the identified features from the environment, either from a single image or a list of images. In Step 3, we collect a dataset from the environment with the VLM-generated extractor code, with which we train an Extractor CNN that will be used in RL training to enable batch computing and increase program speed. The methodology is formalized below.
Step 1: Querying the VLM for Semantic Features. To extract intelligible features from an unseen environment that could be used to train an effective RL policy, we need to first determine what features are useful to model decision-making. In this step, we provide the VLM with:
  • $\mathcal{D}$: Basic domain knowledge about the environment.
  • $\mathcal{I}_{0}$: An example image of the environment.
The VLM identifies key semantic features critical for decision-making, represented as two distinct sets:
  • $\mathcal{F}_{s}=\{f_{1},f_{2},\dots,f_{p}\}$: Features that can derived from a single image frame.
  • $\mathcal{F}_{m}=\{f_{p+1},f_{p+2},\dots,f_{n}\}$: Features requiring multiple frames (e.g., velocity, acceleration, tendency of change). Note that $\mathcal{F}_{m}$ may not be needed in every environment.
The process can be formalized in Eq. 1:
$\mathcal{F}=\mathcal{F}_{s}\cup\mathcal{F}_{m},\quad\mathcal{F}=\mathcal{M}(\mathcal{D},\mathcal{I}_{0}).$ (1)
For example, in the Skiing environment, $\mathcal{F}_{s}$ includes features such as the position of skier, the position of next gate, the distance to next gate, the position of nearest tree, and the distance to nearest tree, while $\mathcal{F}_{m}$ includes the current horizontal speed of the skiier. This example is also included in Fig. 1.
Step 2: Querying the VLM for Feature Extraction Code. In this step, we query the VLM to generate feature extraction code, denoted by $\mathcal{G}$, to retrieve the identified set of features $\mathcal{F}$. To mitigate hallucination, we provide visual segmentations for each image, which will be covered in the discussion section. The visual segmentation $\mathcal{S}$ is computed using a segmentation function, represented as $Seg$, such that $\mathcal{S}=\text{Seg}(\mathcal{I})$. The code generation process involves two stages:
Single Image Extraction: For features in $\mathcal{F}$, the VLM generates code $\mathcal{G}_{s}$ to extract as many features as possible from a single image-segmentation pair $(\mathcal{I},\mathcal{S})$. In practice, as demonstrated in our experiments, this approach effectively recovers all features in $\mathcal{F}_{s}$. To enhance the robustness of the generated code, the VLM is provided with multiple example images during this stage to ensure it has been exposed to all possible scenarios. See Algorithm 1.
Multiple Image Extraction: For features in $\mathcal{F}_{m}$, the VLM generates code $\mathcal{G}_{m}$, which extends $\mathcal{G}_{s}$ to process sequential image-segmentation pairs $\{(I_{t-i},S_{t-i}),\dots,(I_{t},S_{t})\}$. One example of such a pair is provided to the VLM during generation. See Algorithm 2.
The final feature extraction code $\mathcal{G}$ is defined in Eq. 2:
$\mathcal{G}=\begin{cases}\mathcal{G}_{s},&\text{if }\mathcal{F}_{m}=\emptyset;\\ \mathcal{G}_{m}(\mathcal{G}_{s}),&\text{if }\mathcal{F}_{m}\neq\emptyset.\end{cases}$ (2)
Algorithm 1 Generating Extraction Code: Single Image
1:Input: Feature set $\mathcal{F}$, images $\{\mathcal{I}_{i}\}_{i=1}^{n}$, VLM $\mathcal{M}$
2:Initialize $\mathcal{G}_{s}=null$
3:$examples\leftarrow\emptyset$
4:for each image $\mathcal{I}_{i}$ in $\{\mathcal{I}_{i}\}_{i=1}^{n}$ do
5:   $\mathcal{S}_{i}\leftarrow\text{Seg}(\mathcal{I}_{i})$
6:   Add $(\mathcal{I}_{i},\mathcal{S}_{i})$ to $examples$
7:end for
8:$\mathcal{G}_{s}\leftarrow\mathcal{M}_{\text{Generate}}(\mathcal{F},examples)$
Algorithm 2 Generating Extraction Code: Image List
1:Input: Feature set $\mathcal{F}$, images $\{\mathcal{I}_{i}\}_{i=1}^{n}$, $\mathcal{G}_{s}$, VLM $\mathcal{M}$
2:Initialize $\mathcal{G}_{m}=null$
3:$examples\leftarrow\emptyset$
4:for each image $\mathcal{I}_{i}$ in $\{\mathcal{I}_{i}\}_{i=1}^{n}$ do
5:   $\mathcal{S}_{i}\leftarrow\text{Seg}(\mathcal{I}_{i})$
6:   Add $(\mathcal{I}_{i},\mathcal{S}_{i})$ to $examples$
7:end for
8:$\mathcal{F}_{m}\leftarrow{\mathcal{M}_{\text{JudgeMultiFrame}}}(\mathcal{F})$
9:if $\mathcal{F}_{m}\neq\emptyset$ then
10:   $\mathcal{G}_{m}\leftarrow\mathcal{M}_{\text{Generate}}(\mathcal{F},examples,\mathcal{G}_{s})$
11:end if
12:Return: $\mathcal{G}_{m}$
Step 3: Dataset Generation and Training an Extractor CNN. In this step, we execute $\mathcal{G}$ on trajectories collected from the environment to generate the feature dataset. For each trajectory, a series of images $\mathcal{T}=\{I_{1},I_{2},\dots,I_{k}\}$ is generated, with corresponding visual segmentations $\mathcal{S}_{t}=\text{Seg}(I_{t})$ for each $t$. The extractor code $\mathcal{G}$ processes each image-segmentation pair $(I_{t},S_{t})$ independently to extract features $\mathbf{v}_{t}=\{v_{1},v_{2},\dots,v_{p}\}$, where $v_{i}\in\mathbb{R}$.
The dataset, consisting of image-feature pairs $\{(\mathcal{I}_{t},\mathbf{v}_{t})\}$, is then used to train an Extractor CNN as a batch-operable and efficient replacement to $\mathcal{G}_{s}$. The trained model predicts semantic features directly from raw images, enabling significantly faster feature extraction during RL training.
If the features require temporal context (i.e., $\mathcal{F}_{m}\neq\emptyset$), $\mathcal{G}_{m}$, consisting of simple mathematical operations on the outputs of $\mathcal{G}_{s}$, is manually adapted to utilize the Extractor CNN in place of $\mathcal{G}_{s}$. For example, in the Skiing environment, $\mathcal{G}_{m}$ retrieves the static features $\mathcal{F}_{s}$ by directly obtaining the values from $\mathcal{G}_{s}$ applied to the last image and acquires the only temporal feature: skiier horizontal speed through subtracting the skiier positions between the first and last image frames. The manual adaptation preserves the exact same logic while ensuring that the subtraction operation is properly batched, significantly accelerating the training process by several orders of magnitude. Without this adaptation, training would be prohibitively time-consuming.
Discussion. Recognizing that hallucination is a well-documented challenge in generative VLM models [33], we provide segmentation along with raw images in Step 2 and 3, using a pre-defined segmentation function $Seg$ that detects objects and their positions in an image by contour and color analysis (details provided in Appendix VII-B). This strategy aligns with common practices in the field [34, 35]. By incorporating segmentation, we mitigate the risk of hallucination and ensure that the generated features and codes are grounded in accurate visual data.
TABLE I: Prompting statistics for VLM-generated code. “Code failure” indicates code that cannot import or run. “Excluded as constant” refers to code that generates invariant outputs across diverse samples and is therefore excluded. “Majority selection coverage” represents the proportion of code yielding identical outputs in the majority voting process. Step 1 refers to the process of generating $\mathcal{G}_{s}$, while Step 2 refers to the process of generating $\mathcal{G}_{m}$.
Environment Type Code Failure Rate (%) Excluded as Constant (%) Majority Selection Coverage (%)
Cart-pole Step 1 0 0 38
Cart-pole Step 2 0 0 100
Skiing Step 1 28 14 26
Skiing Step 2 0 0 100
Boxing Step 1 20 46 14
Boxing Step 2 - - -
TABLE II: Performance of Extractor CNN in extracting feature vectors across different environments, compared with ground truth from VLM-generated code. Continuous components are evaluated with MSE, normalized to [0,1], and scaled by $10^{-3}$; boolean components are evaluated with accuracy. Dashes indicate vacant entries due to varying feature lengths and types across environments.
Environment Feature Vector Indices (1–11)
1 2 3 4 5 6 7 8 9 10 11
Skiing (MSE) 0.11 0.02 0.16 0.20 1.06 2.13 1.93 0.50 1.81
Cart-pole (MSE) 0.008 0.016
Boxing (MSE) 1.07 2.30 23.25 19.42 6.80 0.98 21.71 20.26 12.12
Boxing (Accuracy) 0.993 0.990
III-B Tree-based RL Training
After obtaining the Extractor CNN, we train an interpretable tree-based model with RL to connect the semantic features with action outputs. During RL training, our framework utilizes the pre-trained Extractor CNN to extract semantic features from image observations, as shown in Fig. 1. The policy network takes these semantic features as input and predicts a distribution over actions, from which an action is sampled and applied to the controlled agent at each time step. Unlike traditional RL algorithms that use black-box models as the policy network, we use an Interpretable Control Tree (ICT) as our policy network to achieve transparent decision-making and human-aligned reasoning while remaining differentiable for end-to-end optimization.
Interpretable Control Tree: Interpretable Control Tree (ICT) is adapted from Interpretable Continuous Control Tree (ICCT) proposed by Paleja et al.  [6]. The ICT is structured similarly as ICCT, where each decision node determines the routing of input features, $\mathbf{x}$, to sparse linear controllers at the leaf nodes.
Each decision node, $i$, is parameterized by weights $\mathbf{w}_{i}$, a bias $b_{i}$, and a steepness parameter $\alpha$ to output a fuzzy Boolean in $[0,1]$, as shown in Eq. 3.
$y_{i}=\sigma\left(\alpha\left(w_{i}^{k}x^{k}-b_{i}\right)\right)$ (3)
The most important feature, $k$, of node, $i$, is selected in Eq. 4 by identifying the weight with the largest magnitude
$k=\arg\max_{j}\left|w_{i}^{j}\right|$ (4)
Since the $\arg\max$ operator is non-differentiable, we approximate this function with the Gumbel-Softmax  [36], as per Eq. 5, where $g_{j}\sim\text{Gumbel}(0,1)$ and $\tau$ is a temperature parameter.
$f(w_{i})^{k}=\frac{\exp\left(\frac{w_{i}^{k}+g_{k}}{\tau}\right)}{\sum_{j}\exp\left(\frac{w_{i}^{j}+g_{j}}{\tau}\right)}$ (5)
During inference, a hard threshold is applied to enforce discrete decision-making using only one feature per node, as per Eq. 6.
$y_{i}=\mathbbm{1}(\alpha(w_{i}^{k}x^{k}-b_{i})>0)$ (6)
While during training, gradients are allowed to flow through discrete decisions using the straight-through trick [37]. This approach ensures that each decision node remains interpretable while allowing for end-to-end gradient-based learning.
Each leaf node, $\ell_{d}$, in the ICT is parameterized by a set of sparse linear controllers, as formulated in Eq. 7, where $\mathbf{\beta}_{d}$ is a sparsely activated weight vector, $\phi_{d}$ is a bias term, and $\mathbf{u}_{d}$ is a $k$-hot encoding vector that enforces sparsity, computed in Eq. 8. The per-leaf selector weight, $\theta_{d}$, learns the relative importance of different input features. The function $h(\cdot)$ provides a differentiable approximation of the $k$-hot function by assigning a value of 1 to the top $k$ elements in $|\mathbf{\theta}_{d}|$ (based on magnitude) and 0 to the rest. Differentiability is preserved using the straight-through trick [37].
$PMF/PDF=(\mathbf{u}_{d}\circ\mathbf{\beta}_{d})^{\top}(\mathbf{u}_{d}\circ\mathbf{x})+\mathbf{u}_{d}^{\top}\phi_{d}$ (7)
$\mathbf{u}_{d}=h\left(\mathrm{softmax}\left(|\mathbf{\theta}_{d}|\right)\right),$ (8)
Unlike ICCT, ICT naturally handles continuous and discrete action space. In the discrete case, leaf nodes in ICT output the possibility of taking each action as a Probability Mass Function (PMF). If there are six possible actions in the discrete action space, then there will be six sparse linear functions of the input features followed by a softmax function at each leaf node. And for the continuous action space, leaf nodes output the Probability Density Function (PDF) over possible actions.
IV Experiments
IV-A Environment
We evaluate our approach on two Atari environments: Skiing and Boxing, and a classic control domain: Cart-pole, where we use image snapshots instead of the built-in feature vector as the observation space. The OpenAI Atari platform is preferable because all their environments provide raw image state spaces without any pre-defined feature vectors or feature sets. This ensures that the VLM engages in genuine reasoning over the environment to identify key semantic features, rather than relying on prior internet-based knowledge, a phenomenon referred to as training data memorization [38].
IV-B Feature Extraction
IV-B1 VLM Deducing Key Semantic Features.
For this step, we utilize GPT-4o, the current state-of-the-art in language reasoning. Across all three environments, the VLM consistently produces a deterministic set of semantic features for decision-making, achieving performance comparable to human reasoning. For all three environments, we query the VLM 50 times and use majority vote to get the most prominent answer from the VLM. We present the identified feature set for the Skiing environment in Fig. 1. Other domains are included in Appendix VII-D.
(a) Boxing
(b) Cartpole
(c) Skiing
Figure 2: Results across three domains. Higher scores are better.
IV-B2 VLMs for Feature Extraction Code
For this step, we utilize claude-3.5-sonnet-20241022, the current state-of-the-art in code generation. To address common challenges with VLM-generated code, such as non-functionality or runtime errors, we query the VLM 50 times using the same prompt (temperature = 0 for reproducibility) and select the most reliable feature extraction code through majority voting.
The generated codes are evaluated on a randomly collected dataset of 100 images and filtered based on the following criteria: (1) the code must compile and execute successfully and (2) it must produce diverse feature outputs rather than constant or repetitive values across the dataset. Candidates producing identical outputs are grouped into clusters, and the largest cluster is chosen as the final feature extraction code. Table I presents statistics from this process, including code failure rates, exclusions due to constant outputs, and majority selection coverage. The majority selection coverage appears low because features related to relative distances are normalized differently, while features related to object positions remain identical.
It is important to note that while Step 1, which generates $\mathcal{G}_{s}$ for feature extraction from a single image frame, is essential for all tasks, Step 2, which generates $\mathcal{G}_{m}$ for feature extraction across multiple image frames, is only necessary when the task’s features depend on temporal information. In the Boxing environment, the feature space does not depend on temporal inference, making Step 2 unnecessary. Consequently, Step 2 is marked as empty for Boxing in Table I. However, the VLM was still queried 50 times, and in all 50 cases, it successfully determined that no code generation was needed, achieving a 100% success rate under this metric.
IV-B3 Extractor CNN Performance Analysis
In Step 3 of our methodology, we address the inefficiencies of the VLM-generated feature extraction code for single-image frames, $\mathcal{G}_{s}$, by training an Extractor CNN as its replacement. While $\mathcal{G}_{s}$ can be executed directly on the dataset, the processing pipeline is inherently non-batchable, as it handles image segmentations independently (see Appendix VII-B for examples). This non-batchability severely impacts RL training, causing slowdowns of 200–2000× compared to batchable operations (Extractor CNN), as demonstrated in Table III. From practical experience, this would translate to training times exceeding multiple weeks on an H40-level GPU for a single policy to converge, making this infeasible.
To constrain computing time to a reasonable range, we train an Extractor CNN to replace $\mathcal{G}_{s}$ and efficiently compute $\mathcal{F}_{s}$, the features derived from a single image frame. The Extractor CNN is specifically designed to resolve the slowdown caused by segmentation-based analysis in $\mathcal{G}_{s}$. For each environment, we train an Extractor CNN on 7,000 images collected from diverse trajectories, validate it on 2,000 images, and test it on 1,000 images. The models achieved a mean squared error (MSE) consistently below 0.02 for each feature in the output vector when compared to the VLM-generated ground truth. Table II provides detailed performance statistics. These results confirm that the Extractor CNN is a reliable and computationally efficient replacement for $\mathcal{G}_{s}$.
TABLE III: Runtime speedup for processing a batch of images across different environments on the same device (averaged over 10 iterations). Results denoted with * are derived from a different device with a better gpu.
Env. Batch VLM (s) CNN (s) Speedup
Skiing 8 0.870 0.011 81.02x
16 1.745 0.003 523.37x
32 3.523 0.005 743.77x
64 7.055 0.008 877.90x
128 14.449 0.021 693.74x
Cart-pole 8 12.726 0.017 738.57x
16 25.679 0.040 637.06x
32 50.204 0.066 763.96x
64 99.660 0.111 898.09x
128* 254.484 0.097 2617.22x
Boxing 8 0.400 0.017 23.07x
16 0.785 0.004 191.03x
32 1.581 0.007 238.51x
64 3.146 0.009 344.34x
128 6.240 0.025 254.40x
IV-C Training Results
Figure 3: Visualization of ICT in the Skiing Environment.
Figure 4: Demonstration of the discrepancy between the physics simulation and the visualization.
We train ICT, our proposed interpretable and verifiable method that learns a tree structure with RL, alongside several baseline models across all environments. Below is an explanation of each baseline model used in the evaluation:
  • CNN: An uninterpretable black-box model that maps raw visual inputs directly to the action space, trained with RL.
  • VLM-ICT: Our proposed interpretable and verifiable method, using an ICT to connect the semantic feature space with the action space, trained with RL.
  • VLM-ICT (Rand Init): Our proposed method, but initialized randomly and not trained with RL.
  • VLM-DT Heuristic: A baseline where VLM generates a heuristic tree to connect the semantic feature space with the action space.
  • VLM-MLP: An uninterpretable benchmark that connects a multi-layer perceptron between the semantic feature space and the action space.
  • Random Policy: A policy that selects actions by randomly sampling from the action space.
Across all three environments, VLM-ICT consistently outperforms other uninterpretable models (MLP) built on the semantic feature space, and demonstrates performance close to the CNN+RL black-box benchmark representing the upper limit of task performance. Notably, in the Cart-pole environment, VLM-ICT even outperforms the black-box CNN model. This is because the CNN fails to accurately capture the pole’s angle, a critical feature for formulating an effective policy. These consistent results highlight the robustness and effectiveness of our proposed method.
The learned ICT in Skiing domain is shown in Fig. 3. This tree was generated fully autonomously using the VLM $\rightarrow$ CNN $\rightarrow$ RL pipeline. The learned trees (ICT) are interpretable both subjectively and objectively, provided they remain small enough, as demonstrated in a previous user study [6]. For the visualization of our learned ICT in Boxing and Cart-pole domains, please refer to Appendix VII-A.
V Limitations and Future Works
In the Cart-pole environment, we observe that the performance of trained VLM-ICT and VLM-MLP has yet to match that of RL policies trained using ground truth features [39, 6]. This discrepancy stems from the gap between the underlying physics simulation and its visual representation, as illustrated in Fig. 4. While the trained Extractor CNN’s predictions closely approximate the feature vectors extracted by the VLM-generated code, they still exhibit deviations from the ground truth feature vectors. This highlights a fundamental challenge in semantic RL from visual observations, where the extracted features may not perfectly align with the underlying physical state. Despite this limitation, our framework still outperforms the CNN-based policy, demonstrating the promise of VLMs in RL. To the best of our knowledge, we are the first to attempt solving the Cart-pole problem using image-based observations, demonstrating a novel and promising direction for interpretable RL.
Future works could explore RL-based backpropagation to fine-tune the VLM models, enabling large language models to iteratively refine their outputs by learning from feedback generated through interactions with RL agents [40]. This approach has the potential to enable VLMs to identify discrepancies between the underlying physics simulation and its visual representation, leading to improved feature extraction for downstream RL tasks.
VI Conclusion
We introduce SILVA, a novel framework for achieving automated semantic interpretability in RL. It automates feature space construction and extraction with zero-shot VLM querying, making it generalizable to unseen environments without human-in-the-loop. To overcome VLMs’ inability to generate valid decision policies, SILVA leverages an ICT for RL training, obtaining transparent, human-understandable, and verifiable decision rules. The framework outperforms black-box models built on the same feature space and achieves performance on par with purely uninterpretable black-box models. While all modern interpretable models, including tree-based approaches, face challenges in complex environments, SILVA stands out among them as the first to achieve automated semantic interpretability in RL, marking a significant advancement that underscores its impact.
Ethics and Impact Statement
SILVA enhances interpretability in RL by automating semantic feature extraction via VLMs, reducing reliance on human annotations and fostering transparent AI decision-making. By structuring policies in an interpretable tree-based format, SILVA aligns with ethical AI principles emphasizing explainability, accountability, and trust.
However, reliance on VLMs introduces biases from pre-trained models, potentially causing discriminatory behavior or unreliable feature extraction. While SILVA improves interpretability, it still depends on black-box models for feature generation, challenging full transparency.
To mitigate these risks, future work could incorporate bias audits, human-in-the-loop verification, and standardized fairness metrics.